Hardwig Inc.
Hardwig Inc. is considering whether to pursue a restricted or relaxed current asset investment policy. The firm's annual sales are expected to total $3,600,000, its fixed assets turnover ratio equals 4.0, and its debt and common equity are each 50% of total assets. EBIT is $150,000, the interest rate on the firm's debt is 10%, and the tax rate is 40%. If the company follows a restricted policy, its total assets turnover will be 2.5. Under a relaxed policy its total assets turnover will be 2.2.

Refer to the data for Hardwig, Inc. Assume now that the company believes that if it adopts a restricted policy, its sales will fall by 15% and EBIT will fall by 10%, but its total assets turnover, debt ratio, interest rate, and tax rate will all remain the same. In this situation, what's the difference between the projected ROEs under the restricted and relaxed policies?

a. 2.46%
b. 3.27%
c. 2.24%
d. 2.98%
e. 2.70%
Answer #1
Restricted Relaxed
Sales 3060000 3600000
Total Assets 1224000 1636363.636
Debt 612000 818181.8182
Equity 612000 818181.8182
EBIT 135000 150000
Less: Interest 61200 81818.18182
EBT 73800 68181.81818
Less: Tax 29520 27272.72727
Net Income 44280 40909.09091
ROE 7.24% 5.00%
Difference 2.24%
Hence, the answer is c.

1 A B С Restricted Relaxed 2 Sales =3600000*85% =3600000 3 Total Assets =B2/2.5 =C2/2.2 4 Debt =B3*50% =C3*50% 5 Equity =B3-B